OpenStudy (anonymous):

10x-4=2x+8 Is the answer -3/-2? solve for x: Need quick help pls

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it would be?

OpenStudy (anonymous):

no it will be 6

OpenStudy (anonymous):

10x-4=2x+8 10x-2x=8+4 8x=12 x=6

OpenStudy (anonymous):

get it ??

OpenStudy (anonymous):

You gotta divide the -8x=-12

OpenStudy (anonymous):

yup divided by (-8) it will be 6